OpenStudy (anonymous):

The sum of two consecutive odd numbers is 44. What is the smaller number?

OpenStudy (anonymous):

21+23=44

OpenStudy (mathstudent55):

Let the smaller number be x The next larger (consecutive) odd numbe is 2 more than x, so it's x + 2 The sum of the numbers x and x + 2 is 44, so x + x + 2 = 44 Solve the equation for x.

OpenStudy (anonymous):

In case you need to do any more problems like these, when you need odd or even consecutive numbers, set up an equation like so: x+(x+2)+(x+4)+...=N In this case, we would have had: x+(x+2)=44 2x+2=44 2x=42 x=21 x is our 1st term, x+2 is our 2nd, and so on. If we just have a problem with consecutive numbers (not odd or even), replace the 2 with a 1, and just continue on adding 1 to each term: x+(x+1)+(x+2)+(x+3)...=N